Grade 12Algebra

Find:
1/7 of 2/9
Find:
1/7 of 2/9
Find:
1/7 of 2/9
Find:
1/7 of 2/9

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (1/7) × (2/9)
By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (1/7) × (2/9)
= (1 × 2)/ (7 × 9)
= (2/63)

Thanks